I need help with this problem
4/3+2/3q=-5/3-1/3q-4

Answer by stanbon(75887)

(4/3)+(2/3)q = (-5/3)-(1/3)q-4
-----------
Multiply thru by 3 to rid of those denominators:
4 + 2q = -5 - q - 12
----
3q = -21
q = -7
================
Cheers,
stan H.